OpenStudy (anonymous):

Please help. I need to solve sec (3theta/2)=-root2. Thank you!

OpenStudy (e.mccormick):

\[\sec \left(\frac{3\theta}{2}\right)=-\sqrt2\] OK, so where are you having trouble with it?

OpenStudy (e.mccormick):